# BAP v2: An Enhanced Task Framework for Instruction Following in Minecraft Dialogues

### v2

@@ 6.3.1 Experiments and Results
Overall trends align with those observed for the GRU-based models in Table 5. The synthetic data, joint training on the aggregated dataset, and CL, all also improve the BERT-based model’s performance on both synthetic and real data, highlighting the robustness of our approach. $M_{\mathrm{agg}}^{\prime}$ achieves an F1 score of 29.4% on $D_{\mathrm{mc}}$, up from $M_{\mathrm{mc}}^{\prime}$ ’s 23.0%, a notable 6.4-percentage point improvement. It is also highly performant across the synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$), surpassing $M_{\mathrm{mc}}^{\prime}$ by large margins on these. On $D_{\mathrm{mc}}$, CL again yields the best-performing model, as with the GRU-based models. $M_{\mathrm{agg+cl}}^{\prime}$ achieves an F1 of 32.3%, a notable 2.9-point improvement over $M_{\mathrm{agg}}^{\prime}$, and an even more significant 9.3-point improvement over $M_{\mathrm{mc}}^{\prime}$. It also achieves significant gains in performance over $M_{\mathrm{agg}}^{\prime}$ on the synthetic data as well.
Comparison to GRU-based models
The lower performance of the BERT-based models than their respective GRU-based counterparts’ is more pronounced for performance on the synthetic data than the Minecraft $D_{\mathrm{mc}}$ data. $M_{\mathrm{mc}}^{\prime}$ achieves an F1 of 23.0% on $D_{\mathrm{mc}}$, 4.3 points lower than $M_{\mathrm{mc}}$’s 27.3%. This gap between the GRU-based and BERT-based counterparts reduces significantly as we move on to look at $M_{\mathrm{agg}}^{\prime}$ and $M_{\mathrm{agg+cl}}^{\prime}$. It goes down to 0.9 and 0.7 points respectively. $M_{\mathrm{agg+cl}}^{\prime}$ is therefore not too far below $M_{\mathrm{agg+cl}}$. The gap in synthetic data performance between $M_{\mathrm{agg}}^{\prime}$ and $M_{\mathrm{agg}}$ is fairly significant. But CL reduces this quite a bit – $M_{\mathrm{agg+cl}}^{\prime}$ is very close to $M_{\mathrm{agg+cl}}$ on $D_{\mathrm{ss}}$, although still behind on the other two. Two key factors contribute to the relatively lower performance of the BERT-based models compared to their GRU counterparts:
  • As mentioned earlier, the sizes of all four datasets were optimized for the GRU-based models. These might need separate tuning for the BERT-based models (and will probably need to be scaled up also as transformers are more complex and data-hungry than GRUs).
  • For tractability of experimentation, we only experimented with the transformer’s and projection layer’s hyperparameters, keeping other architectural hyperparameters identical to the GRU-based models (e.g. those of the CNNs).
These choices were intentional, as our goal here (as stated in the beginning of this section) was primarily to demonstrate the robustness of our synthetic data and training approach, and not to optimize dataset sizes or architectural hyperparameters for maximizing performance. We leave such fine-tuning for future work.
Model Train Data Test Data
$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
$M_{\mathrm{mc}}^{\prime}$ $D_{\mathrm{mc}}$ 29.1 17.8 12.0 23.0
$M_{\mathrm{agg}}^{\prime}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 70.2 55.0 36.8 29.4
$M_{\mathrm{agg+cl}}^{\prime}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 75.9 67.4 45.3 32.3
Table 6: BERT- and transformer-based models
@@ 7 Model Analysis
We now analyze the performance and behavior of the best model $M_{\mathrm{agg+cl}}$ (Section 6.2.2) relative to the baseline $M_{\mathrm{mc}}$ (Section 2.6). In Section 7.1, we do so using the BAP v2 evaluation benchmark introduced in Section 4.4. Section 7.2 complements this with a qualitative evaluation of outputs from both models.
@@ 7.1 Quantitative evaluation
EB Baseline ($M_{\mathrm{mc}}$) 76.4 75.2 53.5 56.0 53.2
Best ($M_{\mathrm{agg+cl}}$) 81.4 80.0 61.8 65.5 61.6
NEB Baseline ($M_{\mathrm{mc}}$) 63.8 60.8 24.6 33.8 24.2
Best ($M_{\mathrm{agg+cl}}$) 66.0 63.2 30.2 39.4 29.6
Overall Baseline ($M_{\mathrm{mc}}$) 65.2 62.4 27.7 36.3 27.3
Best ($M_{\mathrm{agg+cl}}$) 67.7 65.0 33.6 42.2 33.0
Table 7: Micro F1 scores for the baseline and best models under the v2 benchmark
Dataset Model Type Color Location Shape Overall
EB Baseline ($M_{\mathrm{mc}}$) 80.9 79.8 61.5 64.9 61.2
Best ($M_{\mathrm{agg+cl}}$) 83.1 81.7 69.0 72.6 68.6
NEB Baseline ($M_{\mathrm{mc}}$) 73.1 69.9 25.7 39.0 25.2
Best ($M_{\mathrm{agg+cl}}$) 73.7 71.2 32.1 46.5 31.5
Overall Baseline ($M_{\mathrm{mc}}$) 73.8 70.7 28.5 41.1 28.1
Best ($M_{\mathrm{agg+cl}}$) 74.5 72.0 35.0 48.6 34.5
Table 8: Macro F1 scores for the baseline and best models under the v2 benchmark
Tables 7 and 8 present all metrics obtained using the v2 benchmark, augmenting Tables 1 and 2 (Section 4.4) with results for the best model. Overall, the best model outperforms the baseline across all metrics. Key takeaways are discussed below.
We begin with the micro scores in Table 7. On the overall dataset, the best model outperforms the baseline on all three F1 metrics: Type, Color, and Location. As noted in Section 4.4, Location F1 is the primary bottleneck to achieving high performance in this task. The best model achieves an overall F1 of 33%, a notable 5.7 percentage-point improvement over the baseline’s 27.3%. This increase in overall F1 is primarily driven by the improvement in Location F1, which similarly rises from 27.7% to 33.6%. These results validate the motivation and impact of our synthetic data. The best model outperforms the baseline in both EB and NEB scenarios. However, as discussed in Section 4.4 regarding the importance of performance on EB, there remains room for improvement: the best model reaches only 61.6%, in spite of the increase from the baseline’s 53.2%. This insight can inform further error analysis and targeted improvements in future models for this task.
Turning to the macro scores in Table 8, we observe an increase across all metrics compared to the micro scores, but the same trends persist. The best model achieves an overall F1 of 34.5% on the overall test set, a notable 6.4 percentage-point improvement over the baseline’s 28.1%. For Shape F1 on the overall test set, the best model scores 42.2% (micro) and 48.6% (macro), up from the baseline’s 36.3% and 41.1%, marking significant gains of 5.9 and 7.5 percentage points, respectively.
@@ 7.2 Qualitative evaluation
We present examples of output from both the baseline and best models, demonstrating the substantial improvement of the latter. For each example, we also discuss the overall F1 metric and relevant auxiliary metrics. A complete table of scores for all metrics across examples is provided in the appendix (Table 10).
Figure 17: Example 1
Example 1
As shown in Figure 17, copying substructures is a common strategy in the MDC. Here, A instructs B to build an arrowhead, asking B to "mirror" the completed half on the opposite side. The baseline model only captures the "on the top" aspect of this instruction, while the best model fully understands and completes the arrowhead. This example also involves placing a floating block. The best model accomplishes this too, by placing and then removing a supporting block (not shown in the figure). It achieves an overall F1 score of 1.0, compared to the baseline model’s 0. However, both models achieve perfect type and color F1 scores, as the baseline model correctly identifies the need for a single purple block to be placed (in terms of net actions).
Figure 18: Example 2
Example 2
As shown in Figure 18, A gives a complex instruction requiring blocks to meet several highly specific criteria, including a criterion involving the builder’s position and orientation. The baseline model only grasps the "two other orange blocks" aspect but poorly interprets their placement. In contrast, the best model fully understands the instruction. Similar to the previous example, this example also involves placing floating blocks, which the best model accomplishes successfully. It achieves an overall F1 score of 1.0, compared to the baseline model’s 0. However, both models achieve perfect type and color F1 scores, as the baseline model correctly identifies the need for two orange blocks to be placed (in terms of net actions).
Figure 19: Example 3
Example 3
As shown in Figure 19, this example mirrors the previous one but focuses on block removals instead of placements. The initial structure, $S_{H}^{t-1}$, consists of three steps extending from a base square. A instructs B to remove a very specific block, using terms like "center" and "second highest step" to describe its location. The baseline model incorrectly selects the third highest step instead of the second, while the best model interprets the instruction accurately. The latter achieves an overall F1 score of 1.0, compared to the former’s 0. However, both models achieve perfect type and color F1 scores, as the baseline model correctly identifies the need to remove one green block.
8 Discussion
We now discuss four topics in detail: an additional key challenge in BAP evaluation and its implications for future work (Section 8.1), the importance of evaluation on the synthetic data (Section 8.2), relevant concurrent work (Section 8.3), and broader implications of our work that go beyond BAP (Section 8.4).
8.1 An additional key challenge for BAP evaluation
Through manual inspection of random samples during the qualitative analysis of model outputs (Section 7.2), we identified an additional challenge for fair evaluation in the BAP task, separate from the issue addressed by the fairer F1 metric in Section 4.1. Specifically, and rather unsurprisingly, the F1 metric lacks the sensitivity to distinguish nuanced differences in error quality. It is possible for two models to make different types of errors, with one type being qualitatively "better" than the other – yet the F1 metric may not capture this distinction effectively. This limitation can hinder fair evaluation and reduce our ability to distinguish between models based on their F1 scores. Below, we illustrate this challenge through qualitative examples. Across all examples, the best model $M_{agg+cl}$ consistently performs qualitatively better than the baseline $M_{mc}$ and is arguably closer to the human reference. However, the F1 metric lacks the sensitivity to reflect this substantial quality gap. In some cases, the auxiliary metrics proposed in Section 4.2 offer some help. A full table of scores for all metrics across examples is provided in the appendix (Table 11). It’s worth noting that such examples were frequent enough in our random inspection and the ones discussed here are fairly representative of the problem, not cherry-picked to favor the best model.
This analysis underscores the inherent challenges in automated evaluation for the BAP task, highlighting the need for further research in this area. The examples also illustrate the difficulty of creating a clear, unambiguous rubric for grading model output on an absolute scale, making human evaluation equally complex. (In fact, we initially considered human evaluation but ultimately dismissed it due to these challenges.)
Figure 20: Example A
@@ 8.1 An additional key challenge for BAP evaluation
As shown in Figure 20, A instructs B to place a yellow block meeting multiple highly specific criteria, including one involving the builder’s position and orientation. The best model largely understands these criteria and places the correct block, though it also adds an extraneous one on each side. In contrast, the baseline model performs poorly, producing seemingly random placements.
The F1 metric rates the baseline model as 0 and the best model as 0.5, indicating an improvement. However, this score does not fully capture the qualitative gap between the two models, especially in spatial reasoning. While both models produce two false positives, the errors made by the best model demonstrate a clearer understanding of the task, an aspect that the F1 metric fails to reflect adequately.
Figure 21: Example B
@@
As shown in Figure 21, A and B are building a 2D giraffe, and A instructs B to place three orange blocks diagonally to form the neck, opposite the tail. The baseline model places three orange blocks, but arranges them horizontally and closer to the tail. While the best model recognizes the need for a diagonal structure, it places it in the wrong location and orientation.
Both models receive an F1 score of 0, with identical scores of 1, 1, and 0 for type, color, and location F1, respectively. However, shape F1 distinguishes the models more effectively: the baseline achieves a shape F1 of 0.33, while the best model achieves a perfect 1.
Figure 22: Example C
@@
As shown in Figure 22, A instructs the addition of two rows of orange blocks directly below each hanging row on either side of the arch. The baseline model places two orange blocks on top, while the best model attaches blocks to each hanging row—still incorrect but a noticeably closer interpretation of A’s intent.
Here, too, the F1 metric scores both models as 0. Unlike the previous example, however, none of the auxiliary metrics distinguish between the models, as both receive identical scores.
Figure 23: Example D

### v3

@@ 6.3 Broader Takeaways
    By providing the LLM-based models with an input representation that is more comparable to the one used by the GRU-based models, we show that our synthetic data significantly improves performance of both non-LLM- and more modern LLM-based models on both the synthetic and original Minecraft data. This highlights the robustness of our data and simple training methodology across different model architectures.
    Motivating future work
    With the fairer v2 evaluation now possible, we can more accurately gauge current performance on the BAP task and the remaining gap for improvement. While Llama-CRAFTS’s 53.0 F1 score marks improved performance, it also underscores that even with SOTA LLMs, significant room for improvement remains, thus reinforcing the task’s challenging nature and its importance for future research.
    Importance of evaluation on the synthetic data
    Going forward, evaluation on the synthetic data will also be important for making more meaningful progress on the BAP task. First, the more structured and simpler nature of the synthetic data allows us to assess more basic model competencies than is possible on the more complex $D_{\mathrm{mc}}$. Second, as noted above, both GRU and LLM baselines $M_{\mathrm{mc}}$ and $M_{\mathrm{mc}}^{\prime}$ underperform significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$; they also perform lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. This counterintuitive behavior shows that models (including SOTA LLMs) trained solely on $D_{\mathrm{mc}}$ may not be as robust. Evaluating on the synthetic data will help build models that do not show such behavior (like we have shown by using the data effectively for training), and complements evaluation on $D_{\mathrm{mc}}$. In fact, future models should aim to surpass previous performance on $D_{\mathrm{mc}}$ while also achieving strong results on synthetic data, ideally improving on both.
    Model Input Train Data Test Data
    $D_{\mathrm{sb}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
    $M_{\mathrm{mc}}^{\prime}$ (Nebula) NN $D_{\mathrm{mc}}$ 49.5 31.9 33.3 47.0
    $M_{\mathrm{agg}}^{\prime}$ NN $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 75.5 54.9 43.8 47.8
    $M_{\mathrm{mc+posb}}^{\prime}$ $N+\textsc{Pos}_{\textbf{B}}$ $D_{\mathrm{mc}}$ 49.7 32.2 32.0 47.6
    $M_{\mathrm{agg+posb}}^{\prime}$ $N+\textsc{Pos}_{\textbf{B}}$ $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 93.4 84.9 82.3 52.5
    $M_{\mathrm{agg+posb+s}}^{\prime}$ (Llama-CRAFTS) $N+\textsc{Pos}_{\textbf{B}}+S$ $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 92.3 83.7 82.2 53.0
    Table 6: Llama performance with different datasets and input representations.
    6.4 Related Work: Role of specialized datasets and smaller finetuned models
    These results highlight a key contemporary question in applied NLP: the relative merits of fine-tuning smaller, specialized models versus leveraging large, general-purpose models via ICL. The performance parity between Nebula and GPT-4 on this task (Section 2.7) suggests that a smaller, fine-tuned specialist model can be as effective as a much larger generalist one. This finding coupled with the fact that we are able to improve the former by a substantial 12.8% via leveraging synthetic data, motivates a deeper exploration of the fine-tuning paradigm and the role of specialized datasets, especially concerning the most effective pathways to further improve performance.
    While generalist models offer impressive few-shot capabilities, avenues for their improvement often center on prompt engineering, increased test-time computation, or scaling to even larger (and often proprietary) models, which may not always be practical. In contrast, the fine-tuning approach offers more vectors for enhancement, e.g., synthetic data (like we show), better training regimes, etc. in addition to the ones for ICL. Furthermore, smaller specialized and open models present practical advantages in terms of computational efficiency and accessibility for both academic research and deployment in resource-constrained settings. For instance, in real-world counterparts of BAP such as instruction-following robots, stringent low-latency requirements often render the extensive test-time computation demanded by larger models infeasible. This pragmatic consideration is supported by emerging theoretical work suggesting that fine-tuning is not merely a matter of convenience. Studies indicate that fine-tuned models can achieve generalization capabilities comparable to ICL (Mosbach et al., 2023; Zhao et al., 2025; Ponce and Etchegoyhen, 2025) and may learn representations that are more efficiently embedded in lower-dimensional manifolds, potentially better capturing real-world data/features (Janapati and Ji, 2025).
    Finally, the development of specialized datasets need not be an endeavor exclusive to fine-tuning. The synthetic data we have generated, rich in specific linguistic and spatial phenomena, could also potentially enhance ICL performance. For instance, it could be used to select more effective few-shot exemplars, drawing inspiration from works like Kulkarni et al. (2024), or be integrated into curriculum learning frameworks for ICL, as explored by Liu et al. (2024), given that the synthetic data is simpler than the original BAP data. This suggests that creating high-quality, task-specific data is a valuable pursuit that can advance multiple modeling paradigms.
@@ 7 Model Analysis
    We now analyze the performance of the best LLM-based model Llama-CRAFTS ($M_{\mathrm{agg+posb+s}}^{\prime}$) relative to the LLM-based baseline Nebula ($M_{\mathrm{mc}}^{\prime}$). In Section 7.1, we do so using the BAP v2 evaluation benchmark introduced in Section 4.3. Section 7.2 complements this with a qualitative evaluation of outputs from both models.
@@ 7.1 Quantitative evaluation
    EB Nebula 92.6 92.0 74.7 80.9 74.2
    Llama-CRAFTS 94.7 94.7 80.3 85.8 80.3
    NEB Nebula 78.8 77.6 44.3 57.1 43.5
    Llama-CRAFTS 80.8 79.5 50.4 59.4 49.4
    Overall Nebula 80.4 79.2 47.7 59.8 47.0
    Llama-CRAFTS 82.4 81.2 53.8 62.4 53.0
    Table 7: F1 scores for the LLM-based baseline Nebula and Llama-CRAFTS
    As can be seen from Table 7, Llama-CRAFTS outperforms Nebula across all metrics on the v2 benchmark, reaching an overall F1 of 53.0, a notable 6.0 points improvement over Nebula ’s 47.0. This increase is primarily driven by an improvement in Location F1, which similarly rises from 47.7 to 53.8. However, the BAP task remains far from being solved. Although the empty board (EB) can be seen as the most basic variant of the BAP task (Section 4.3), the best model’s 80.3 (albeit a substantial increase from the baseline’s 74.2) is a clear indication that there remains much room for improvement.
@@ 7.2 Qualitative evaluation
    Scenario 1: One model is correct, the other makes mistakes
@@
    Figure 11: The overall F1 score is sufficient to distinguish between correct and incorrect model outputs (Scenario 1)
    Scenario 2: The auxiliary metrics provide further insight